OpenStudy (anonymous):

d=vt+1/2at^2 solve for t

OpenStudy (anonymous):

\[ \left(\frac 12 a\right)t^2+(v)t+(-d) = 0 \]

OpenStudy (anonymous):

Are you familiar with the quadratic equation?

OpenStudy (anonymous):

quadratic formula^

OpenStudy (anonymous):

http://www.purplemath.com/modules/quadform.htm In this case the variable is \(t\) instead of \(x\).